OpenStudy (anonymous):

looking to integrate a differential equation on both sides... d^2C/dz^2 = -zA* dC/dz how would I go about doing this?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

thanks! i'll be sure to give credit where credit is due. just to let you know, I simplified the constants in the equation into "A"

OpenStudy (amistre64):

you replied to an automated message that really does nothing but post that you have aquestion in the chatroom box :)

OpenStudy (amistre64):

there you are :)

OpenStudy (anonymous):

thanks I noticed that when i clicked around. need an answer though!

OpenStudy (amistre64):

d^2C/dz^2 = -zA* dC/dz hmm.... havent done to many of these so I dont know alot of the techniques, except for swapping variables when doing a first derivative..

OpenStudy (anonymous):

its a secondary order differential equation.

OpenStudy (amistre64):

C'' = -zA C' ...

OpenStudy (amistre64):

Id say divide both sides by C'... but thats just a guess :)

OpenStudy (amistre64):

its outta my league.... good luck with it :)

OpenStudy (anonymous):

d^2C/dz^2 = -zA* dC/dz (D^2+Az)C=0 where D=d/dz

OpenStudy (anonymous):

thanks anyway. its for transport phenomenon in biological systems :P

OpenStudy (anonymous):

solution says that form is in a form of a natural exponential. can anyone explain?

OpenStudy (anonymous):

i tried to recall the solution of higher order DE, but could not,away from it for a long time,i just remember the formation of auxilliary eq